OpenStudy (anonymous):

on the parallel line, find the ordered pair where x=-2 still using y-3=3(x+1)

OpenStudy (the_fizicx99):

Do you have the slope?

OpenStudy (anonymous):

I believe it's 3

OpenStudy (anonymous):

4x I belive.

OpenStudy (the_fizicx99):

The slope is M so your saying its: y - 3 = 3(x-1) ?

OpenStudy (the_fizicx99):

@XbTst4U how did you get 4x?

OpenStudy (anonymous):

are we talking about the slope of "y-3=3(x+1)" or the slope of x=-2?

OpenStudy (the_fizicx99):

x can't have a slope.. I'm asking if you have the slope for the equation you've asked.

OpenStudy (compassionate):

Hello, Parallel lines have the same slope. So you just want to put the equation in y= mx + b x=-2 still using y-3=(x+1) Okay, so solve the equation first. y = 3x + 3 Now, we know x = -2 y = 3(-2) + 3 y = -6 + 3 y = -3 The line is horizontal and has a slope of 0.

OpenStudy (anonymous):

I see the problem the equation is y-3=3(x+1) some how my pc dropped a number compassionate is the slope still o

OpenStudy (the_fizicx99):

Then its just, y-3=3x + 3 +3 +3 y = 3x + 6

OpenStudy (anonymous):

y=3x=6 x=(-4) or (-4,0) the equation in standard form of a parallel line that passes though (0,-2) is -3x+y=-2

OpenStudy (anonymous):

I found the answer if any one need it y-3=3(x+1) y=3x+3+3 y=3x+6 y-(-2=m(x-0) m=3 y+2=3x -3x+y+2=0 3x-y-2=0